# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ IV.2.3 Visualizing the coupling effects
In Fig. 17B, the impact of coupling on oscillatory dynamics between two FHN nodes is shown, highlighting a tendency towards antiphase synchronization. To quantify this synchronization, we employ a methodology capable of determining the characteristic synchronization time applicable across various oscillatory scenarios (Fig. 17C). This approach involves analyzing the correlation across individual oscillations and tracking the evolution of this correlation over successive cycles, aligning with techniques described by Toral et al. Toral et al. (2003a). Our findings also reveal that the number of cycles required to reach the ultimate asynchronous state varies with the coupling strength (Fig. 17D).
Despite the model’s simplicity, it encapsulates a broad spectrum of dynamics Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005a); Stefanescu and Jirsa (2008); Yanagita et al. (2005), setting the stage for exploring more complex systems involving multiple oscillators and connections, often used in the study of chimera states.
@@ IV.3 Chimera states
Chimera states, known for their intriguing spatiotemporal patterns characterized by coexisting domains of coherence and incoherence, have garnered significant attention in studies of coupled FHN oscillators Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Sawicki et al. (2019); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Chouzouris et al. (2018); Schmidt et al. (2017). This dynamical regime can manifest across diverse network configurations. Work by Omelchenko et al. Omelchenko et al. (2013, 2015) offers valuable insights into the underlying mechanisms and conditions conducive to chimera states.
@@
$\varepsilon{u_{k}}_{t}=-\frac{u_{k}^{3}}{3}+u_{k}-v_{k}+\\ \frac{\sigma}{2R}\sum_{k-R}^{k+R}[b_{uu}(u_{j}-u_{k})+b_{uv}(v_{j}-v_{k})]$ (39a)
${v_{k}}_{t}=u_{k}+a_{k}+\frac{\sigma}{2R}\sum_{k-R}^{k+R}[b_{vu}(u_{j}-u_{k})+b_{vv}(v_{j}-v_{k})]$ (39b)
with $R$ representing the radius of interaction among adjacent nodes. Following Omelchenko et al. (2013, 2015) we chose $b=0$, $\varepsilon=0.05$ and $\sigma=0.1$. Furthermore, the coupling terms of the equation are defined by a coupling rotation matrix,
$\begin{pmatrix}b_{uu}&b_{uv}\\ b_{vi}&b_{vv}\end{pmatrix}=\begin{pmatrix}\cos{\phi}&\sin{\phi}\\ -\sin{\phi}&\cos{\phi}\end{pmatrix},$ (40)
@@
Figure 18 illustrates the manifestation of chimera states. The analysis often involves mean phase velocities, indicative of the system’s dynamical state. In these states, a segment of incoherently oscillating nodes coexists with a synchronized group, forming distinct coherent and incoherent domains in the network. This phenomenon aligns with findings from Omelchenko et al. Omelchenko et al. (2015), where increasing initial variation gradually merges the two distinct regions.
@@
space
In summary, the study of coupled oscillators unveils a vast range of dynamic behaviors, attributed to the diverse network structures, coupling mechanisms, and the inherent parameter versatility within these systems. For more straightforward configurations, analytical insights into the system’s bifurcation and stability aspects can be obtained Hoff et al. (2014); Campbell and Waite (2001). The application of coupled oscillator models extends across various fields, notably in modeling neuronal Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020), electrical Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006); Takembo et al. (2019); Ciszak et al. (2003), and biological systems Kori et al. (2012); Vodeneev et al. (2011); Sukhov et al. (2011); Cartwright (2000). A significant portion of research focuses on understanding synchronization patterns across nodes Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Sawicki et al. (2019); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Chouzouris et al. (2018); Schmidt et al. (2017), often employing correlation functions for analysis Toral et al. (2003a). Additionally, the role of noise in influencing system dynamics garners considerable attention, revealing its impact on the coupled oscillators’ behavior Toral et al. (2003a); Yu et al. (2021); Lindner (2004); Neiman et al. (1999); Brandstetter et al. (2010); Semenova et al. (2016); Perc (2007, 2005a); Zaks et al. (2005); García-Ojalvo and Schimansky-Geier (1999).
@@ V Conclusions
We structured our analysis into three primary sections. Firstly, we examined the original FHN model [Eq. (8)], discussing widely observed dynamical regimes such as monostability, multistability, relaxation oscillations, and excitability. We examined the role of local and global bifurcations in shaping these regimes, emphasizing the importance of time scale separation. Secondly, we explored the diffusively coupled FHN model [Eq. (10)], introducing spatial coupling through diffusion. Through theoretical analysis, we investigated stationary homogeneous solutions, their linear stability, and spatially structured dynamic solutions, including traveling structures and spatially extended patterns. We studied the emergence of a Turing instability and the resulting spatially structured Turing patterns. Additionally, we examined front solutions, localized states, traveling pulses, and pacemaker-driven waves within the oscillatory domain, highlighting the richness of patterns that arose in different spatial dimensions. Lastly, we explored discretely coupled FHN equations [Eq. (11)]. This is the broadest category as here one can consider a multitude of different network topologies and coupling terms. We focussed ourselves on synchronization properties in two coupled FHN modules, the existence of traveling waves when transitioning from continuous diffusive coupling to discrete coupling, and the emergence of chimera states characterized by spatiotemporal patterns of coherent and incoherent behavior.
In conclusion, our review serves as a guide for understanding and using the diverse dynamical behaviors offered by the FHN model. Throughout our analysis, stability analyses and bifurcation studies provided insights into the observed dynamics, offering a comprehensive understanding of the FHN model’s capabilities and applications. By exploring its applications across multiple disciplines, we aimed to inspire further exploration and application of the FHN model in diverse scientific domains.
@@ VI Discussion
The dynamics of the FHN model are influenced by various factors, including external driving, internal feedback, inter-system coupling, noise, and time delays. These elements contribute not only to the model’s inherent complexity but also to its extensive applicability across physical and biological domains. While our review has examined in detail the dynamics of single and coupled FHN systems, it has not extensively covered the roles of external driving, self-feedback, time delays, and noise. Each of these aspects warrants further exploration for a comprehensive understanding of the FHN model’s dynamics and potential. Living systems are typically non-equilibrium, active, and multi-scale complex systems that often operate in transient states away from stable attractors. The FHN model is an excellent candidate to serve as a testbed for developing new methods to analyze such transient dynamics and non-reciprocal interactions. Integrating the FHN model within multiphysics frameworks, which encapsulate diverse physical interactions, is essential for advancing our understanding of these systems. Such new techniques and integrative models are particularly important in the complex simulation of biological phenomena. Cardiac and neuronal examples, in particular, highlight the FHN model’s adaptability and significance in these areas.
External Driving.: Introducing an external stimulus to the FHN system can induce resonance or forced oscillations, when the stimulus frequency aligns with the system’s inherent frequency. This aspect is pivotal in understanding biological rhythms and has applications in electronic and mechanical systems, as well as in neuroscience for modeling brain rhythms and treating disorders like Parkinson’s. Researchers have looked into the dynamics of externally driven FHN models, revealing their potential for controlling excitable behaviors Ullner et al. (2003a); Pankratova et al. (2005); Kuramoto et al. (1998). Of course, the external driving could also come from another FHN system, leading to two autonomous dynamical systems in a master–slave configuration. Synchronisation has been studied in such unidirectionally coupled models, also in the presence of additional external aperiodic forcing Ciszak et al. (2003).
@@
Time delays. Time delays play an important role in nonlinear systems, extensively researched due to their pervasive presence across a spectrum of physical and biological contexts. These delays, which can be intentionally introduced via feedback loops or inherently present in various systems, play a significant role in shaping the behavior of dynamical models. In systems ranging from optical and electronic devices to neuronal networks and technological infrastructures, time delays emerge due to the finite speed of signal transmission, processing times, and the effects of memory and latency. Incorporating time delays into models offers new strategies for controlling and designing nonlinear systems, enhancing their stability and performance. In neural networks, for instance, the inherent time delays in signal transmission, owing to diverse neural pathways, introduce complex dynamics that can be replicated in models like the FHN framework, not only through delayed interactions but as an intrinsic characteristic of the system. These delays can significantly impact the network’s behavior, leading to phenomena such as enhanced synchronization, the suppression of undesirable oscillations, and the formation of chimera states Schülen et al. (2019); Perlikowski et al. (2010); Saha and Feudel (2017); Achouri et al. (2022); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Jia et al. (2015); Chun-Biao et al. (2010); Toral et al. (2003b); Yu et al. (2021).
@@
Neuronal function, for instance, is inherently noisy, influenced by the stochastic opening and closing of ion channels, variable presynaptic currents, and conductivity fluctuations. Lindner et al. used the FHN model as as prototype of excitable stochastic dynamics to investigate the effects of Gaussian white noise on such systems. Their research spanned isolated units to networks of coupled elements, unveiling phenomena like noise-induced oscillations, stochastic resonance, and synchronization, alongside noise-triggered phase transitions and complex pulse and spiral dynamics, with applications extending from biophysics to laser technology Lindner (2004).
Counterintuitively, random perturbations, especially when combined with weak deterministic stimuli, can introduce order in temporal and spatial domains of nonlinear systems, a phenomenon known as stochastic resonance. This effect, where noise enhances system response to external perturbations, underscores noise’s constructive potential in nonlinear dynamics. Coherence resonance occurs when an excitable system exhibits maximal signal regularity at optimal noise levels, even without external driving forces. This phenomenon, alongside stochastic resonance, has been pivotal in understanding dynamics near bifurcation points and within bistable and oscillatory systems. Also in the context of the FHN model, it has sparked significant interest for its implications in synchronization Yu et al. (2021); Neiman et al. (1999); Brandstetter et al. (2010); Makarov et al. (2001), resonance Lindner and Schimansky-Geier (1999); Toral et al. (2003a); Lindner (2004); Wiesenfeld et al. (1994); Longtin (1993); Perc (2005b); Brandstetter et al. (2010); Semenova et al. (2016); Zaks et al. (2005); Pankratova et al. (2005); Sagués et al. (2007), the formation and stability of complex spatio-temporal patterns like spirals Perc (2007); Toral et al. (2003a); García-Ojalvo and Schimansky-Geier (1999); Sagués et al. (2007); Ullner et al. (2003b); García-Ojalvo et al. (2001).
Transient dynamics. Many techniques developed in nonlinear dynamics focus on analyzing the stability and structure of attracting solutions. However, living systems, in particular, are most often in transient Hastings et al. (2018); Rabinovich et al. (2008); Verd et al. (2014) and continuously integrate time-varying, noisy information from both internal and external sources. Currently, there is a lack of a clear framework to characterize such long transient dynamics of system state trajectories. The need to adapt the mathematical formalism when describing biological systems in changing environments has recently been emphasized Koch and Koseska (2024); Koch et al. (2024); Nandan and Koseska (2023).
@@
Neuronal models. The brain’s complexity necessitates a multiscale modeling approach, from single neuron dynamics to network behaviors Deco et al. (2008); Siettos and Starke (2016). The FHN model provides a straightforward yet potent representation of neuron dynamics and their integration into larger neuronal models. Bridging the microscopic and macroscopic scales is key to understanding brain function. Mathematical models, from biophysical to data-driven, and tools from statistics to dynamical systems, help decipher the brain’s intricate structure and functionality. These models facilitate hypothesis testing within biological contexts, shedding light on brain connectivity and dynamics from neurons to networks.
@@ Acknowledgements.
The numerical codes to reproduce the figures in this study are openly available in GitLab Git , and as an archived repository on RDR by KU Leuven [Upcoming].

### v4

@@ IV.2.3 Visualizing the coupling effects
In Fig. 17B, the impact of coupling on oscillatory dynamics between two FHN nodes is shown, highlighting a tendency towards antiphase synchronization. To quantify this synchronization, we determine the characteristic synchronization time for different synchronization strengths (Fig. 17C). This approach involves analyzing the correlation across individual oscillations and tracking the evolution of this correlation over successive cycles, aligning with techniques described by Toral et al. Toral et al. (2003a). This reveals that the number of cycles required to reach the ultimate antiphase synchronized state varies with the coupling strength (Fig. 17D).
Despite the model’s simplicity, it encapsulates a broad spectrum of dynamics Vaidyanathan (2015); Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Keener (1988); Lindner et al. (2004); Neiman et al. (1999); Heinrich et al. (2010); Perc (2005b); Stefanescu and Jirsa (2008); Yanagita et al. (2005), setting the stage for exploring more complex systems involving multiple oscillators and connections, often used in the study of chimera states.
@@ IV.3 Chimera states
Chimera states, known for their intriguing spatio-temporal patterns characterized by coexisting domains of coherence and incoherence, have garnered significant attention in studies of coupled FHN oscillators Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Sawicki et al. (2019); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Chouzouris et al. (2018); Schmidt et al. (2017); Huang et al. (2024b). This dynamical regime can manifest across diverse network configurations. Work by Omelchenko et al. Omelchenko et al. (2013, 2015) offers valuable insights into the underlying mechanisms and conditions conducive to chimera states.
@@
$\begin{array}[]{ll}\varepsilon{u_{k}}_{t}=-\frac{u_{k}^{3}}{3}+u_{k}-v_{k}+\frac{\sigma}{2R}\sum_{k-R}^{k+R}[b_{uu}(u_{j}-u_{k})+b_{uv}(v_{j}-v_{k})],\vspace{3pt}\\ {v_{k}}_{t}=u_{k}+a_{k}+\frac{\sigma}{2R}\sum_{k-R}^{k+R}[b_{vu}(u_{j}-u_{k})+b_{vv}(v_{j}-v_{k})],\end{array}$ (35)
with $R$ representing the radius of interaction among adjacent nodes. Following Omelchenko et al. (2013, 2015) we chose $b=0$, $\varepsilon=0.05$ and $\sigma=0.1$. Furthermore, the coupling terms of the equation are defined by a rotation matrix,
$\begin{pmatrix}b_{uu}&b_{uv}\\ b_{vu}&b_{vv}\end{pmatrix}=\begin{pmatrix}\cos{\phi}&\sin{\phi}\\ -\sin{\phi}&\cos{\phi}\end{pmatrix},$ (36)
@@
Fig. 18 illustrates the manifestation of chimera states. The analysis often involves mean phase velocities, indicative of the system’s dynamical state. In these states, a segment of incoherently oscillating nodes coexists with a synchronized group, forming distinct coherent and incoherent domains in the network. This phenomenon aligns with findings from Omelchenko et al. Omelchenko et al. (2015), where increasing initial variation gradually merges the two distinct regions.
@@
In summary, the study of coupled oscillators unveils a vast range of dynamic behaviors, attributed to the diverse network structures, coupling mechanisms, and the inherent parameter versatility within these systems. For simpler configurations, analytical insights into the system’s bifurcation and stability aspects can be obtained Hoff et al. (2014); Campbell and Waite (2001). The application of coupled oscillator models extends across various fields, in modeling neuronal Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020), electrical Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006); Takembo et al. (2019); Ciszak et al. (2003), and biological systems Kori et al. (2012); Vodeneev et al. (2011); Sukhov et al. (2011); Cartwright (2000); Cebrián-Lacasa et al. . A significant portion of research focuses on understanding synchronization patterns across nodes Schöll (2016a); Omelchenko et al. (2013); Nikitin et al. (2019); Sawicki et al. (2019); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Chouzouris et al. (2018); Schmidt et al. (2017), often employing correlation functions for analysis Toral et al. (2003a). Additionally, the role of noise in influencing system dynamics receives considerable attention, revealing its impact on the coupled oscillators’ behavior Toral et al. (2003a); Yu et al. (2021); Lindner et al. (2004); Neiman et al. (1999); Brandstetter et al. (2010); Semenova et al. (2016); Perc (2007, 2005b); Zaks et al. (2005); García-Ojalvo and Schimansky-Geier (1999).
@@ V Conclusions
We structured our analysis into three primary sections. Firstly, we examined the original FHN model [Eq. 8], discussing widely observed dynamical regimes such as monostability, multistability, relaxation oscillations, and excitability. We examined the role of local and global bifurcations in shaping these regimes, emphasizing the importance of time scale separation. Secondly, we explored the diffusively coupled FHN model [Eq. 10], introducing spatial coupling through diffusion. Through theoretical analysis, we investigated stationary homogeneous solutions, their linear stability, and spatially structured dynamic solutions, including traveling structures and spatially extended patterns. We studied the emergence of a Turing instability and the resulting spatially structured Turing patterns. Additionally, we examined front solutions, localized states, traveling pulses, and pacemaker-driven waves within the oscillatory domain, highlighting the richness of patterns that arose in different spatial dimensions. Lastly, we explored discretely coupled FHN equations [Eq. 11]. This is the broadest category as here one can consider a multitude of different network topologies and coupling terms. We focussed on synchronization properties in two coupled FHN modules, the existence of traveling waves when transitioning from continuous diffusive coupling to discrete coupling, and the emergence of chimera states characterized by spatio-temporal patterns of coherent and incoherent behavior.
In conclusion, we hope our review will serve as a guide for understanding and using the diverse dynamical behaviors offered by the FHN model. Throughout our analysis, stability analyses and bifurcation studies provided insights into the observed dynamics. By exploring its applications across multiple disciplines, we aimed to inspire further exploration and application of the FHN model in diverse scientific domains.
@@ VI Discussion
The dynamics of the FHN model are influenced by various factors, including external driving, internal feedback, inter-system coupling, noise, and time delays. These elements contribute not only to the model’s inherent complexity but also to its extensive applicability across physical and biological domains. While our review has examined the dynamics of single and coupled FHN systems in some detail, it has not extensively covered the roles of external driving, self-feedback, time delays, and noise. Each of these aspects warrants further exploration for a comprehensive understanding of the FHN model’s dynamics and potential. Living systems are typically non-equilibrium, active, and multi-scale complex systems that often operate in transient states away from stable attractors. The FHN model is an excellent candidate to serve as a testbed for developing new methods to analyze such transient dynamics and non-reciprocal interactions. Integrating the FHN model within multiphysics frameworks, which encapsulate diverse physical interactions, is essential for advancing our understanding of these systems. Such new techniques and integrative models are particularly important in the complex simulation of biological phenomena. Cardiac and neuronal examples, in particular, highlight the FHN model’s adaptability and significance in these areas.
External Driving. Introducing an external stimulus to the FHN system can induce resonance or forced oscillations, when the stimulus frequency aligns with the system’s inherent frequency. This aspect is pivotal in understanding biological rhythms and has applications in electronic and mechanical systems, as well as in neuroscience for modeling brain rhythms and treating disorders like Parkinson’s. Researchers have looked into the dynamics of externally driven FHN models, revealing their potential for controlling excitable behaviors Ullner et al. (2003a); Pankratova et al. (2005); Kuramoto et al. (1998). Of course, the external driving could also come from another FHN system, leading to two autonomous dynamical systems in a master-slave configuration. Synchronization has been studied in such unidirectionally coupled models, also in the presence of additional external aperiodic forcing Ciszak et al. (2003).
@@
Time delays. Time delays play an important role in nonlinear systems, extensively researched due to their pervasive presence across a spectrum of physical and biological contexts. These delays, which can be intentionally introduced via feedback loops or inherently present in various systems, play a significant role in shaping the behavior of dynamical models. In systems ranging from optical and electronic devices to neuronal networks and technological infrastructures, time delays emerge due to the finite speed of signal transmission, processing times, and the effects of memory and latency. Incorporating time delays into models offers new strategies for controlling and designing nonlinear systems, enhancing their stability and performance. In neural networks, for instance, the inherent time delays in signal transmission, owing to diverse neural pathways, introduce complex dynamics that can be replicated in models like the FHN framework, not only through delayed interactions but as an intrinsic characteristic of the system. These delays can significantly impact the network’s behavior, leading to phenomena such as enhanced synchronization, the suppression of undesirable oscillations, and the formation of chimera states Schülen et al. (2019); Perlikowski et al. (2010); Saha and Feudel (2017); Achouri et al. (2022); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Jia et al. (2015); Chun-Biao et al. (2010); Toral et al. (2003b); Yu et al. (2021); Elfouly et al. (2024).
@@
Neuronal function, for instance, is inherently noisy, influenced by the stochastic opening and closing of ion channels, variable presynaptic currents, and conductivity fluctuations. Lindner et al. used the FHN model as prototype of excitable stochastic dynamics to investigate the effects of Gaussian white noise on such systems Lindner et al. (2004). Their research spanned isolated units to networks of coupled elements, unveiling phenomena like noise-induced oscillations, stochastic resonance, and synchronization, alongside noise-triggered phase transitions and complex pulse and spiral dynamics, with applications extending from biophysics to laser technology Lindner et al. (2004).
Counterintuitively, random perturbations, especially when combined with weak deterministic stimuli, can introduce order in temporal and spatial domains of nonlinear systems, a phenomenon known as stochastic resonance. This effect, where noise enhances system response to external perturbations, underscores noise’s constructive potential in nonlinear dynamics. Coherence resonance occurs when an excitable system exhibits maximal signal regularity at optimal noise levels, even without external driving forces. This phenomenon, alongside stochastic resonance, has been pivotal in understanding dynamics near bifurcation points and within bistable and oscillatory systems. Also in the context of the FHN model, it has sparked significant interest for its implications in synchronization Yu et al. (2021); Neiman et al. (1999); Brandstetter et al. (2010); Makarov et al. (2001), resonance Lindner and Schimansky-Geier (1999); Toral et al. (2003a); Lindner et al. (2004); Wiesenfeld et al. (1994); Longtin (1993); Perc (2005a); Brandstetter et al. (2010); Semenova et al. (2016); Zaks et al. (2005); Pankratova et al. (2005); Sagués et al. (2007), the formation and stability of complex spatio-temporal patterns like spirals Perc (2007); Toral et al. (2003a); García-Ojalvo and Schimansky-Geier (1999); Sagués et al. (2007); Ullner et al. (2003b); García-Ojalvo et al. (2001).
Transient dynamics. Many techniques developed in nonlinear dynamics focus on analyzing the stability and structure of attracting solutions. However, living systems, in particular, are most often in transient Hastings et al. (2018); Rabinovich et al. (2008); Verd et al. (2014) and continuously integrate time-varying, noisy information from both internal and external sources. Currently, there is a lack of a clear framework to characterize such long transient dynamics of system state trajectories. The need to adapt the mathematical formalism when describing biological systems in changing environments has recently been emphasized Koch and Koseska (2024); Koch et al. (2024); Nandan (2023).
@@
Neuronal models. The brain’s complexity requires a multiscale modeling approach, from single neuron dynamics to network behaviors Deco et al. (2008); Siettos and Starke (2016). The FHN model provides a straightforward yet potent representation of neuron dynamics and their integration into larger neuronal models. Bridging the microscopic and macroscopic scales is key to understanding brain function. Mathematical models, from biophysical to data-driven, and tools from statistics to dynamical systems, help decipher the brain’s intricate structure and functionality. These models facilitate hypothesis testing within biological contexts, shedding light on brain connectivity and dynamics from neurons to networks.
@@ Acknowledgements.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT in order to receive suggestions on grammar and phrasing. All scientific content and figures were created by the authors. After using this tool, the authors reviewed the suggestions and only edited the grammar and phrasing of parts of the text. The authors take full responsibility for the content of the publication.
@@
The numerical codes to reproduce the figures in this study are openly available in GitLab Git (2024), and as an archived repository on RDR by KU Leuven Cebrián-Lacasa et al. (2024).